OpenStudy (anonymous):

For a game, Tony has a rectangle drawn on a piece of paper that has an area of 18 in.2. What should he do to the dimensions in order to have a similar rectangle that's area is only 2 in.2? Divide them by nine. Divide them by six. Divide them by three. Multiply them by three.

OpenStudy (anonymous):

unit analysis area of small rectangle = 1/9 * area of large rectangle (m^2) length of small rectangle = x * length of large rectangle (m) length^2 of small rectangle = x^2 * length^2 of large rectangle (m^2) x^2=1/9 x=?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

x=1/3; divide by 3